# Dual-Channel Multiplex Graph Neural Networks for Recommendation

### v1

@@ Abstract.
Efficient recommender systems play a crucial role in accurately capturing user and item attributes that mirror individual preferences. Some existing recommendation techniques have started to shift their focus towards modeling various types of interaction relations between users and items in real-world recommendation scenarios, such as clicks, marking favorites, and purchases on online shopping platforms. Nevertheless, these approaches still grapple with two significant shortcomings: (1) Insufficient modeling and exploitation of the impact of various behavior patterns formed by multiplex relations between users and items on representation learning, and (2) ignoring the effect of different relations in the behavior patterns on the target relation in recommender system scenarios. In this study, we introduce a novel recommendation framework, Dual-Channel Multiplex Graph Neural Network (DCMGNN), which addresses the aforementioned challenges. It incorporates an explicit behavior pattern representation learner to capture the behavior patterns composed of multiplex user-item interaction relations, and includes a relation chain representation learning and a relation chain-aware encoder to discover the impact of various auxiliary relations on the target relation, the dependencies between different relations, and mine the appropriate order of relations in a behavior pattern. Extensive experiments on three real-world datasets demonstrate that our DCMGNN surpasses various state-of-the-art recommendation methods. It outperforms the best baselines by 10.06% and 12.15% on average across all datasets in terms of $R@10$ and $N@10$ respectively. The source code of our paper is available at https://anonymous.4open.science/r/KDD24-5FDE.
@@ 3. Preliminary
A basic behavior pattern between users $\mathcal{U}$ and items $\mathcal{V}$ in multiplex bipartite graphs is defined as $\mathcal{U}\xrightarrow{{[{r_{1}}]\&[{r_{2}}]\&\ldots\&[{r_{|\mathcal{R}|}}]}}\mathcal{V}$ that represents a unique interaction behavior pattern between users and items, where []delimited-[]normal-⋅[$\cdot$][ ⋅ ] denotes optional, $|\mathcal{R}|$ is the number of relation types, and at least one interaction relation type $r_{i}$ exists.
@@ 5. Experiments
Datasets
Users
Items
Interactions
Relation Type
Retail
2,174
30,113
$9.7\times 10^{4}$
{View, Cart, Buy}
Tmall
15,449
11,953
$1.2\times 10^{6}$
{View, Cart, Buy}
Yelp
19,800
22,734
$1.4\times 10^{6}$
{Tips, Like, Neutral, Dislike}
@@ A.1. Notations
A.2. Detailed Dataset Description
Three datasets are adopted for evaluation:
@@ A.3. Baselines
A.4. Additional Experiments
A.4.1. Performance on Mitigating Data Sparsity

### v3

@@ Abstract.
Efficient recommender systems play a crucial role in accurately capturing user and item attributes that mirror individual preferences. Some existing recommendation techniques have started to shift their focus towards modeling various types of interaction relations between users and items in real-world recommendation scenarios, such as clicks, marking favorites, and purchases on online shopping platforms. Nevertheless, these approaches still grapple with two significant shortcomings: (1) Insufficient modeling and exploitation of the impact of various behavior patterns formed by multiplex relations between users and items on representation learning, and (2) ignoring the effect of different relations in the behavior patterns on the target relation in recommender system scenarios. In this study, we introduce a novel recommendation framework, Dual-Channel Multiplex Graph Neural Network (DCMGNN), which addresses the aforementioned challenges. It incorporates an explicit behavior pattern representation learner to capture the behavior patterns composed of multiplex user-item interaction relations, and includes a relation chain representation learning and a relation chain-aware encoder to discover the impact of various auxiliary relations on the target relation, the dependencies between different relations, and mine the appropriate order of relations in a behavior pattern. Extensive experiments on three real-world datasets demonstrate that our DCMGNN surpasses various state-of-the-art recommendation methods. It outperforms the best baselines by 10.06% and 12.15% on average across all datasets in terms of $R@10$ and $N@10$ respectively.
@@ 3. Preliminary
A basic behavior pattern between users $\mathcal{U}$ and items $\mathcal{V}$ in multiplex bipartite graphs is defined as $\mathcal{U}\xrightarrow{{[{r_{1}}]\&[{r_{2}}]\&\ldots\&[{r_{|\mathcal{R}|}}]}}\mathcal{V}$ that represents a unique interaction behavior pattern between users and items, where []delimited-[][$\cdot$][ ⋅ ] denotes optional, $|\mathcal{R}|$ is the number of relation types, and at least one interaction relation type $r_{i}$ exists.
@@ 5. Experiments
Datasets Users Items Interactions Relation Type
Retail 2,174 30,113 $9.7\times 10^{4}$View, Cart, Buy}
Tmall 15,449 11,953 $1.2\times 10^{6}$View, Cart, Buy}
Yelp 19,800 22,734 $1.4\times 10^{6}$Tips, Like, Neutral, Dislike}
@@ A.1. Notations
A.2. Detailed Dataset Description
Three datasets are adopted for evaluation:
@@ A.3. Baselines
A.4. Detailed Experimental Setting
For the recommendation task, we treat the connected user-item pairs as positive node pairs, and consider other unlinked node pairs as negative pairs. We divide the positive node pairs into the training set and test set according to the proportion of 75% and 25%. At the same time we randomly select the same number of negative node pairs to add to the training set and test set.
For our DCMGNN, we set the number of aggregation layers in LightGCN to 2, and $\tau$ to 0.1, then we set $\mu_{1}$ to 0.1 and $\mu_{2}$ to 0.5. For NTMR, we xxxxxx. For MBGCN, we xxxxxx. For MB-HGCN and MB-CGCN, we xxxxxx. For CRGCN, we xxxxxx. For MBCMN, we xxxxxx. For HMG-CR, we xxxxxx. For S-MBRec, we xxxxxx. For CML, we xxxxxx. For KMCLR, we xxxxxx. For BPHGNN, we xxxxxx.
A.4.1. Time and Space Complexity Analysis
The DCMGNN consists of three critical components: explicit behavior pattern representation learner, implicit relation chain effect learner and relation chain-aware contrastive learning.
For explicit behavior pattern representation learner, the time complexity of aggregating all basic behavior pattern is $O(\mathbb{N}N^{2})$, and with the graph convolution, the time complexity of local behavior pattern aggregation is $O(\mathbb{N}N^{2})+O(N^{2}dl+Nd+Nd^{2}(l-1))=O(n^{2}(\mathbb{N}+dl)+Nd+Nd^{2}(l-1))$. Similarly, the time complexity of local behavior pattern aggregation is also $O(n^{2}(\mathbb{N}+dl)+Nd+Nd^{2}(l-1))$. So the total time complexity of explicit behavior pattern representation learner is $O(n^{2}(\mathbb{N}+dl)+Nd+Nd^{2}(l-1))$. For implicit relation chain effect learner, the time complexity of multiplex relation embedding aggregation is $O(N^{2}dl+Nd+Nd^{2}(l-1))$, the time complex of relation chain representation learning is $O(N^{2}dl+Nd)$, so the total time complexity of implicit relation chain effect learner is $O(N^{2}dl+Nd+Nd^{2}(l-1))$. For relation chain-aware contrastive learning, the time complexity is $O(N^{2}d+Nd)$. Therefore, the total time complexity of our DCMGNN is $O((\mathbb{N}+dl)N^{2}+(d+d^{2}(l-1))N)$.
A.5. Additional Experiments
A.5.1. Performance on Mitigating Data Sparsity